# Luther-Emery liquid and dominant singlet superconductivity in the hole-doped Haldane spin-1 chain

### v2

@@ I Introduction
Figure 2: Comparison of correlation functions. The decays of the normalized singlet pair-pair $P_{\mathrm{singlet}}(R)$, density-density $N(R)$, and spin-spin correlations $S(R)$ with distance $R$ are contrasted for (a) $J_{\mathrm{H}}=0.200U$ and (b) $J_{\mathrm{H}}=0.275U$. Both panels are for $U/W=1.6$ and $L=96$ at $x=1/12$ hole doping. The dotted lines indicate a power law decay with exponent $\alpha=1$. The two panels showcase the trend where the singlet pair-pair correlations become dominant and long range (i.e. decaying slower than $R^{-1}$) at high $J_{\mathrm{H}}/U$ values. An expanded version of this figure showing the evolution for additional values of $J_{\mathrm{H}}/U$ is provided in Fig. 2 of Appendix A.
@@ II Model
where $n_{i\gamma\sigma}=c^{\dagger}_{i\gamma\sigma}c^{\vphantom{\dagger}}_{i\gamma\sigma}$ is the number operator, $U>0$ is the intra-orbital Hubbard repulsion, and the second term describes inter-orbital density-density interactions. $J_{\mathrm{H}}$ represents the Hund’s coupling strength. Physically, it is expected that $U^{\prime}>J_{\mathrm{H}}$ [51, 52], which holds for $J_{\mathrm{H}}/U<1/3$. In this paper we report results for $0.2\leq J_{\mathrm{H}}/U\leq 0.35$, where the value $0.35$ is included to show that the results do not change drastically at the boundary value $1/3$. The third term is the Hund’s coupling term, and the fourth is the on-site inter-orbital electron-pair hopping with $P_{i\gamma^{\prime}}=c_{i\gamma^{\prime}\uparrow}c_{i\gamma^{\prime}\downarrow}$. The spin-$1/2$ operators in Eq. (2) are defined as $S_{i\gamma}^{\alpha}=\frac{1}{2}\sum_{\sigma\sigma^{\prime}}c_{i\gamma\sigma}^{\dagger}\tau_{\sigma\sigma^{\prime}}^{\alpha}c^{\vphantom{\dagger}}_{i\gamma\sigma^{\prime}},$ where $\alpha\in\{x,y,z\}$ and $\vec{\tau}=\left(\sigma^{x},\sigma^{y},\sigma^{z}\right)$ is the vector of Pauli matrices.
@@ IV.2 Energy gaps and entanglement
Three signs point towards the possibility of a Luther-Emery liquid state in the hole-doped TOHC: i) there is a finite spin gap $\Delta E_{\mathrm{s}}(2,x)$, ii) long-range singlet pair-pair correlations are observed at large Hund coupling, and iii) the TOHC is formally similar to the two-leg one-band Hubbard ladder at weak coupling, which is considered an archetypal Luther-Emery system. Indeed, the two orbitals can be represented as fictious legs in a two-leg ladder [Fig. 1(b)].
@@ V Conclusion
To realize this physics in materials, two nearly degenerate orbitals are required. This rules out many compounds already known to realize Haldane spin chain physics, such as the nickel-based Y2BaNiO5 [67], which has significant level splitting and may be in an entirely different regime [26]. Nevertheless, quasi-one-dimensional materials with two highly degenerate orbitals are certainly possible, as evidenced by materials such as OsCl4 [68]. However, its $U/W$ ratio may be too high to be relevant for our superconducting mechanism at intermediate coupling, instead justifying a spin-1 chain description [69, 70]. The currently leading candidates are compounds like RuOCl2 and OsOCl2 [71, 72], which have $U\approx W$ and $J_{\mathrm{H}}/U=0.2$. These strongly anisotropic van der Waals materials also feature subleading interchain hoppings within the plane and very weak interplane hoppings [71]. When the purely one-dimensional superconducting state discussed in this article is dominant, such interchain couplings may stablize it into a true long-range order. Further research into candidate materials and experimental realizations of the ideas presented in this publication should be pursued.
@@ Appendix B Additional entanglement properties
Ψgit clone https://github.com/g1257/dmrgpp.git
Ψgit clone https://github.com/g1257/PsimagLite.git
@@
Ψcd PsimagLite/lib; perl configure.pl; make
Ψcd ../../dmrgpp/src; perl configure.pl; make
@@
Ψexport PATH="<PATH-TO-DMRG++>/src:\$PATH"
@@
Ψobserve -f input.ain "<gs|c?0’*c?0;c?1’*c?1|gs>" -p 12 > nupndown.txt
@@
Ψ##Ainur1.0
ΨTotalNumberOfSites=\$N;
ΨNumberOfTerms=5;
ΨModel=HubbardOneBandExtendedSuper;
Ψ
ΨSolverOptions="twositedmrg,calcAndPrintEntropies";
ΨVersion="version";
Ψ
Ψ# Keep a maximum of m states, but allow truncation with tolerance and minimum states
ΨTruncationTolerance="1e-6,100";
ΨInfiniteLoopKeptStates=100;
ΨFiniteLoops=[
Ψ[@auto, 100, 0],
Ψ[@auto, 250, 0],
Ψ[@auto, 500, @save]ΨΨΨΨΨΨ# save output to .hd5 file that can be read by observe
Ψ];
Ψ
Ψ# Set filling and magnetization sector.
ΨTargetElectronsUp=\$L;
ΨTargetElectronsDown=\$L;
Ψ# To search all symmetry sectors, use TargetElectronsTotal=\$N instead.
Ψ
Ψ# Tolerance for Lanczos
Ψstring LanczosOptions="reortho";ΨΨ# enables explicit reorthogonalization
ΨLanczosEps=1e-10;
Ψint LanczosSteps=400;
Ψ
ΨhubbardU=[6.4,...x\$N];
ΨpotentialV=[0,...x\$(2N)];
Ψ
Ψ### Hoppings.
Ψgt0:DegreesOfFreedom=1;
Ψgt0:GeometryKind="ladder";
Ψgt0:LadderLeg=2;
Ψgt0:GeometryOptions="ConstantValues";
Ψgt0:dir0:Connectors=[-1];ΨΨΨΨ# hopping along ladder leg
Ψgt0:dir1:Connectors=[0.0];ΨΨΨΨ# hopping along ladder rung
Ψ
Ψ### U’-JH/2
Ψgt1:DegreesOfFreedom=1;
Ψgt1:GeometryKind="ladder";
Ψgt1:LadderLeg=2;
Ψgt1:GeometryOptions="ConstantValues";
Ψgt1:dir0:Connectors=[0.0];
Ψgt1:dir1:Connectors=[2.0];
Ψ
Ψ### U[2] ~ -2JH. This is the SxSx+SySy part of the Hund’s coupling term.
Ψgt2:DegreesOfFreedom=1;
Ψgt2:GeometryKind="ladder";
Ψgt2:LadderLeg=2;
Ψgt2:GeometryOptions="ConstantValues";
Ψgt2:dir0:Connectors=[0.0];
Ψgt2:dir1:Connectors=[-3.52];
Ψ
Ψ### U[3] ~ -2JH. This is the SzSz part of the Hund’s coupling term.
Ψgt3:DegreesOfFreedom=1;
Ψgt3:GeometryKind="ladder";
Ψgt3:LadderLeg=2;
Ψgt3:GeometryOptions="ConstantValues";
Ψgt3:dir0:Connectors=[0.0];
Ψgt3:dir1:Connectors=[-3.52];
Ψ
Ψ### U[4] ~ JH. This is the pair-hopping term. In this DMRG++ model the sign is positive.
Ψgt4:DegreesOfFreedom=1;
Ψgt4:GeometryKind="ladder";
Ψgt4:LadderLeg=2;
Ψgt4:GeometryOptions="ConstantValues";
Ψgt4:dir0:Connectors=[0.0];
Ψgt4:dir1:Connectors=[1.76];

### v3

@@ I Introduction
Figure 2: Comparison of correlation functions. The decays of the normalized singlet pair-pair $P_{\mathrm{singlet}}(R)$, density-density $N(R)$, and spin-spin correlations $S(R)$ with distance $R$ are contrasted for (a) $J_{\mathrm{H}}=0.200U$ and (b) $J_{\mathrm{H}}=0.275U$. Both panels are for $U/W=1.6$ and $L=96$ at $x=1/12$ hole doping. The dotted lines indicate a power law decay with exponent $\alpha=1$. The two panels showcase the trend where the singlet pair-pair correlations become dominant and long range (i.e. decaying slower than $R^{-1}$) at high $J_{\mathrm{H}}/U$ values. An expanded version of this figure showing the evolution for additional values of $J_{\mathrm{H}}/U$ is provided in Appendix A.
@@ II Model
where $n_{i\gamma\sigma}=c^{\dagger}_{i\gamma\sigma}c^{\vphantom{\dagger}}_{i\gamma\sigma}$ is the number operator, $U>0$ is the intra-orbital Hubbard repulsion, and the second term describes inter-orbital density-density interactions. $J_{\mathrm{H}}$ represents the Hund’s coupling strength. We assume the standard relation $U^{\prime}=U-2J_{H}$, which arises due to spin-rotational invariance. Physically, it is expected that $U^{\prime}>J_{\mathrm{H}}$ [51, 52], which holds for $J_{\mathrm{H}}/U<1/3$. In this paper we report results for $0.2\leq J_{\mathrm{H}}/U\leq 0.35$, where the value $0.35$ is included to show that the results do not change drastically at the boundary value $1/3$. The third term is the Hund’s coupling term, and the fourth is the on-site inter-orbital electron-pair hopping with $P_{i\gamma^{\prime}}=c_{i\gamma^{\prime}\uparrow}c_{i\gamma^{\prime}\downarrow}$. The spin-$1/2$ operators in Eq. (2) are defined as $S_{i\gamma}^{\alpha}=\frac{1}{2}\sum_{\sigma\sigma^{\prime}}c_{i\gamma\sigma}^{\dagger}\tau_{\sigma\sigma^{\prime}}^{\alpha}c^{\vphantom{\dagger}}_{i\gamma\sigma^{\prime}},$ where $\alpha\in\{x,y,z\}$ and $\vec{\tau}=\left(\sigma^{x},\sigma^{y},\sigma^{z}\right)$ is the vector of Pauli matrices.
@@ IV.2 Energy gaps and entanglement
Three signs point towards the possibility of a Luther-Emery liquid state in the hole-doped TOHC: i) there is a finite spin gap $\Delta E_{\mathrm{s}}(2,x)$, ii) long-range singlet pair-pair correlations are observed at large Hund’s coupling, and iii) the TOHC is formally similar to the two-leg one-band Hubbard ladder at weak coupling, which is considered an archetypal Luther-Emery system. Indeed, the two orbitals can be represented as fictious legs in a two-leg ladder [Fig. 1(b)].
@@ V Conclusion
To realize this physics in materials, two nearly degenerate orbitals are required  [^1] Our preliminary results on short chains with fixed $J_{H}/U=0.25$ suggest that i) the binding and pair-pair correlations are robust to the introduction of interorbital hopping, although the optimal $U/W$ ratio changes, ii) the binding and pair-pair correlations survive up to at least 10% detuning of the diagonal intraorbital hopping, and iii) the physics is sensitive to crystal field effects. Indeed, a crystal field raising the energy of one orbital by $0.03W$ already inhibits binding in $L=16$ chains. Thus deviations from a diagonal nearest-neighbor hopping matrix may be acceptable, but care should be taken to find materials with minimal crystal field splitting.. This rules out many compounds already known to realize Haldane spin chain physics, such as the nickel-based Y2BaNiO5 [68], which has significant level splitting and may be in an entirely different regime [26]. Nevertheless, quasi-one-dimensional materials with two highly degenerate orbitals are certainly possible, as evidenced by materials such as OsCl4 [69]. However, its $U/W$ ratio may be too high to be relevant for our superconducting mechanism at intermediate coupling, instead justifying a spin-1 chain description [70, 71]. The currently leading candidates are compounds like RuOCl2 and OsOCl2 [72, 73], which have $U\approx W$ and $J_{\mathrm{H}}/U=0.2$. These strongly anisotropic van der Waals materials also feature subleading interchain hoppings within the plane and very weak interplane hoppings [72]. When the purely one-dimensional superconducting state discussed in this article is dominant, such interchain couplings may stablize it into a true long-range order. Further research into candidate materials and experimental realizations of the ideas presented in this publication should be pursued.
Access to the computational results reported in this paper will be made available from Ref. [74].
@@ Appendix B Additional entanglement properties
git clone https://github.com/g1257/dmrgpp.git
git clone https://github.com/g1257/PsimagLite.git
@@
cd PsimagLite/lib; perl configure.pl; make
cd ../../dmrgpp/src; perl configure.pl; make
@@
export PATH="<PATH-TO-DMRG++>/src:\$PATH"
@@
observe -f input.ain "<gs|c?0’*c?0;c?1’*c?1|gs>" -p 12 > nupndown.txt
@@
##Ainur1.0
TotalNumberOfSites=\$N;
NumberOfTerms=5;
Model=HubbardOneBandExtendedSuper;
SolverOptions="twositedmrg,calcAndPrintEntropies";
Version="version";
# Keep a maximum of m states, but allow truncation with tolerance and minimum states
TruncationTolerance="1e-6,100";
InfiniteLoopKeptStates=100;
FiniteLoops=[
[@auto, 100, 0],
[@auto, 250, 0],
[@auto, 500, @save]ΨΨΨΨΨΨ# save output to .hd5 file that can be read by observe
# Set filling and magnetization sector.
TargetElectronsUp=\$L;
TargetElectronsDown=\$L;
# To search all symmetry sectors, use TargetElectronsTotal=\$N instead.
# Tolerance for Lanczos
string LanczosOptions="reortho";ΨΨ# enables explicit reorthogonalization
LanczosEps=1e-10;
int LanczosSteps=400;
hubbardU=[6.4,...x\$N];
potentialV=[0,...x\$(2N)];
### Hoppings.
gt0:DegreesOfFreedom=1;
gt0:GeometryKind="ladder";
gt0:LadderLeg=2;
gt0:GeometryOptions="ConstantValues";
gt0:dir0:Connectors=[-1];ΨΨΨΨ# hopping along ladder leg
gt0:dir1:Connectors=[0.0];ΨΨΨΨ# hopping along ladder rung
### U’-JH/2
gt1:DegreesOfFreedom=1;
gt1:GeometryKind="ladder";
gt1:LadderLeg=2;
gt1:GeometryOptions="ConstantValues";
gt1:dir0:Connectors=[0.0];
gt1:dir1:Connectors=[2.0];
### U[2] ~ -2JH. This is the SxSx+SySy part of the Hund’s coupling term.
gt2:DegreesOfFreedom=1;
gt2:GeometryKind="ladder";
gt2:LadderLeg=2;
gt2:GeometryOptions="ConstantValues";
gt2:dir0:Connectors=[0.0];
gt2:dir1:Connectors=[-3.52];
### U[3] ~ -2JH. This is the SzSz part of the Hund’s coupling term.
gt3:DegreesOfFreedom=1;
gt3:GeometryKind="ladder";
gt3:LadderLeg=2;
gt3:GeometryOptions="ConstantValues";
gt3:dir0:Connectors=[0.0];
gt3:dir1:Connectors=[-3.52];
### U[4] ~ JH. This is the pair-hopping term. In this DMRG++ model the sign is positive.
gt4:DegreesOfFreedom=1;
gt4:GeometryKind="ladder";
gt4:LadderLeg=2;
gt4:GeometryOptions="ConstantValues";
gt4:dir0:Connectors=[0.0];
gt4:dir1:Connectors=[1.76];